# Measuring and Analyzing Intelligence via Contextual Uncertainty in Large Language Models using Information-Theoretic Metrics

### v2

@@ 6.3 Monotonicity of the Marginal Entropy ($H_{k}$)
Models trained on natural language typically show a steady decline in $H_{k}$, but this is not an information-theoretic necessity. The trend reflects the non-stationary character of language.
@@
We have $H_{k}=H(p(Y_{k+1}))$. If $p(Y_{j})$ were the same for every position $j$, then $H_{k}$ would not change with $k$; a process with position-independent statistics is stationary. Natural language is not stationary: word probabilities differ by position (e.g., first words differ from tenth words). Language models learn this and therefore satisfy $p(Y_{k+1})\neq p(Y_{j+1})$ in general. Hence $H_{k}$ must change with $k$. The observed decline of $H_{k}$ means that as context grows the set of plausible next tokens shrinks, so the marginal distribution becomes more predictable and its entropy falls.
A Bayesian View of the Entropy Metrics.
We can read the three entropy-based quantities as follows:
  • Prior uncertainty: $H_{k}=H(Y_{k+1})$ is the uncertainty before seeing a specific context $X_{1:k}$.
  • Expected posterior uncertainty: $h_{k}=\mathbb{E}[H(Y_{k+1}\mid X_{1:k})]$ is the expected uncertainty after observing the context.
  • Expected information gain: $I_{k}:=I(X_{1:k};Y_{k+1})$ is the expected drop in uncertainty, i.e. the mutual information.
They satisfy the identity
$H_{k}\;=\;h_{k}+I_{k}.$
Proposition 2 (Monotonicity of $H_{k}$ under Predictive Convergence).
If an autoregressive model obeys the Axiom of Predictive Convergence
$h_{k+1}+I_{k+1}\;\leq\;h_{k}+I_{k}\qquad\text{for all }k\geq 0,$
then $H_{k+1}\leq H_{k}$.
Proof.
The identity $H_{j}=h_{j}+I_{j}$ makes the axiom equivalent to $H_{k+1}\leq H_{k}$. ∎
Interpretation.
The axiom says that the model’s total predictive uncertainty cannot grow as it sees more data. Rewriting it as $I_{k+1}-I_{k}\leq h_{k}-h_{k+1}$ yields the Axiom of Diminishing Information Efficiency: the extra mutual information from one more token never exceeds the drop in expected posterior uncertainty.
@@ 6.4 Relationship Between the Decay Rates of $h_{k}$ and $H_{k}$
Define single-step drops $\Delta h_{k}:=h_{k}-h_{k+1}\geq 0$ and $\Delta H_{k}:=H_{k}-H_{k+1}\geq 0$.
@@
The ratio of $\Delta h_{k}$ to $\Delta H_{k}$ shows what kind of information the new token provides:
@@
6.5 A Symmetric View via Bayesian Decomposition
Proposition 3 (Bayesian Decomposition of Marginal Entropy).
For every $k\geq 1$,
$H_{k}-h_{k}\;=\;H(X_{1:k})-H(X_{1:k}\mid Y_{k+1}).$
@@ 6.5 A Symmetric View via Bayesian Decomposition
Mutual information symmetry gives
$H(A)-H(A\mid B)\;=\;H(B)-H(B\mid A).$
Let $A:=Y_{k+1}$ and $B:=X_{1:k}$. Then $H_{k}=H(Y_{k+1})$ and $h_{k}=H(Y_{k+1}\mid X_{1:k})$, yielding the stated identity. ∎
@@
$H_{k}-h_{k}=I(X_{1:k};Y_{k+1})$ measures forward predictive power, while the right-hand side measures how much the next token tells us about the past. They are equal because mutual information is symmetric.
@@ 6.6 Theoretical Analysis of the Information-Gain Span
$\operatorname{IGS}(k_{\mathrm{s}},k_{\mathrm{l}}):=u_{k_{\mathrm{s}}}\bigl(1-u_{k_{\mathrm{l}}}\bigr),\hskip 17.00024ptu_{k}:=h_{k}/H_{k}.$
Theorem 1 (IGS Peak and Markov Order).
Assume the text source is an ergodic Markov chain of exact order $m$ and the model predicts perfectly. Assume further that $H_{k}$ is non-increasing. Then $\operatorname{IGS}(k_{\mathrm{s}},k_{\mathrm{l}})$ is maximised when
$k_{\mathrm{s}}\leq m<k_{\mathrm{l}}.$
@@
Step 1: Behaviour of $u_{k}$. For $k<m$, $h_{k}$ falls strictly and $H_{k}$ does not rise, so $u_{k}$ falls. For $k\geq m$, $h_{k}$ is constant ($h_{m}$) while $H_{k}$ still falls, so $u_{k}$ rises or stays flat. Thus $u_{k}$ has a single minimum at $k=m$.
Step 2: The factor $(1-u_{k_{\mathrm{l}}})$. This term is largest when $u_{k_{\mathrm{l}}}$ is smallest, i.e. when $k_{\mathrm{l}}$ is just beyond $m$ (say $m+1$).
Step 3: The factor $u_{k_{\mathrm{s}}}$. Because $u_{k}$ falls on $k<m$, $u_{k_{\mathrm{s}}}$ is largest when $k_{\mathrm{s}}$ is as small as allowed, subject to $k_{\mathrm{s}}\leq m$.
The product is therefore maximised when $k_{\mathrm{s}}$ is small ($\leq m$) and $k_{\mathrm{l}}$ is just above $m$. ∎
@@
The theorem shows that if a model truly learns a Markov order $m$, then the IGS curve peaks where the short window lies inside that horizon and the long window lies just outside it. By scanning $k_{\mathrm{s}}$ and $k_{\mathrm{l}}$, we can locate the peak and treat its position as an estimate of the model’s effective memory.
@@ 7 Conclusion
We introduced a framework for characterising the information–processing behaviour of large language models (LLMs) via their Entropy Decay Curves (EDCs), obtained by plotting the length-conditional uncertainty index $u_{k}$ against context length $k$. This approach yields dynamic cognitive profiles that go beyond conventional single-value metrics.
Empirically, these profiles vary markedly with (i) model architecture, (ii) parameter scale, and (iii) the intrinsic predictability of the input text. The distinct decay trajectories traced by Llama 3.3, DeepSeek-R1, and Qwen 2.5 reveal different strategies for exploiting contextual information. We further defined the Information Gain Span (IGS), a concise statistic that summarises the overall shape of an EDC and enables direct quantitative comparison between models.
The method also functions as a diagnostic tool. The “entropy collapse” observed for Llama 3.3 on the Alice corpus provides strong quantitative evidence consistent with memorisation arising from pre-training data contamination.
The metrics rest on solid theory. First, we proved that the average conditional entropy $h_{k}$ decreases monotonically under a Predictive Attenuation Condition. Second, we showed that the IGS peaks precisely when additional context no longer improves predictive accuracy.
Future work will apply this methodology to a wider range of models and specialised corpora (e.g., code, medical, legal) to chart how cognitive profiles vary with training domain, pushing toward a deeper, mechanistic understanding of artificial intelligence.

### v3

@@ 5.3 Monotonicity of the Marginal Entropy ($H_{k}$)
Models trained on natural language often exhibit a decline in $H_{k}$ as $k$ grows, but this is not an information-theoretic necessity: $H_{k}$ is an entropy of an average predictive distribution, and mixture effects can in principle make it increase or stay flat.
@@
Recall our theoretical definition from the previous subsection:
$\bar{p}_{k}(y):=\mathbb{E}\!\left[p\!\left(y\mid X^{(k)}\right)\right],\qquad H_{k}:=-\sum_{y\in\mathcal{Y}}\bar{p}_{k}(y)\,\log_{2}\bar{p}_{k}(y).$
Thus $H_{k}$ can depend on $k$ through two mechanisms.
  1. 1.
    Context-length effect. Increasing $k$ changes the conditioning information in $p(y\mid x)$, which can systematically reshape the averaged distribution $\bar{p}_{k}$.
  2. 2.
    Sampling-position effect. In a finite evaluated segment, the sliding-window rule defines $X^{(k)}=(T_{I},\dots,T_{I+k-1})$ with targets at positions $I+k$. Changing $k$ shifts the set of target positions being averaged over, so the induced distribution of contexts $X^{(k)}$ (and hence $\bar{p}_{k}$) may change even if the underlying text is non-stationary.
Our blockwise positional robustness checks (Section 3.4) indicate that, within the tested Alice segment, the estimates of $h_{k}$, $H_{k}$, and $u_{k}$ remain stable across widely separated blocks (for both $k=3$ and $k=600$ in our updated table), suggesting that the sampling-position effect is small in that setting.
Decomposition of Predictive Uncertainty.
Under the sliding-window sampling scheme, a randomly sampled context $X^{(k)}$ induces next-token probabilities $p(y\mid X^{(k)})$ for $y\in\mathcal{Y}$. Averaging these predictive distributions over the sampling rule yields the averaged predictive distribution
$\bar{p}_{k}(y):=\mathbb{E}\!\left[p\!\left(y\mid X^{(k)}\right)\right],\qquad y\in\mathcal{Y},$
whose entropy is
$H_{k}:=-\sum_{y\in\mathcal{Y}}\bar{p}_{k}(y)\,\log_{2}\bar{p}_{k}(y).$
Conversely, the quantity
$h_{k}:=\mathbb{E}\!\left[-\sum_{y\in\mathcal{Y}}p\!\left(y\mid X^{(k)}\right)\log_{2}p\!\left(y\mid X^{(k)}\right)\right]$
is the expected predictive entropy within a randomly sampled context.
We define the information-gain term
$I_{k}:=H_{k}-h_{k},$
which measures the expected reduction in uncertainty when predicting with the context-specific distribution $p(\cdot\mid X^{(k)})$ rather than with the averaged distribution $\bar{p}_{k}$. By Jensen’s inequality (concavity of entropy), $H_{k}\geq h_{k}$, hence $I_{k}\geq 0$, and the decomposition $H_{k}=h_{k}+I_{k}$ holds by definition.
@@ 5.4 Relationship Between the Decay Rates of $h_{k}$ and $H_{k}$
Define single-step differences $\Delta h_{k}:=h_{k}-h_{k+1}$ and $\Delta H_{k}:=H_{k}-H_{k+1}$. Under Proposition 1 we typically have $\Delta h_{k}\geq 0$, whereas $\Delta H_{k}$ need not be nonnegative in general (though it is often observed to be so in practice).
@@
When $\Delta H_{k}>0$, comparing $\Delta h_{k}$ and $\Delta H_{k}$ is a useful diagnostic:
@@
5.5 A Symmetric View via Mutual Information
Proposition 2 (Symmetric decomposition of $H_{k}-h_{k}$).
Fix $k$ and let $X^{(k)}$ be the random context under the sliding-window sampling rule as in the previous subsection. Define an auxiliary next-token random variable $\tilde{Y}^{(k)}$ by
$\Pr(\tilde{Y}^{(k)}=y\mid X^{(k)}=x):=p(y\mid x),\qquad y\in\mathcal{Y}.$
Then the marginal law of $\tilde{Y}^{(k)}$ is $\Pr(\tilde{Y}^{(k)}=y)=\bar{p}_{k}(y)$, hence
$H_{k}=H(\tilde{Y}^{(k)}),\qquad h_{k}=H(\tilde{Y}^{(k)}\mid X^{(k)}).$
Consequently,
$H_{k}-h_{k}=I\!\left(X^{(k)};\tilde{Y}^{(k)}\right)=H\!\left(X^{(k)}\right)-H\!\left(X^{(k)}\mid\tilde{Y}^{(k)}\right).$
@@ 5.5 A Symmetric View via Mutual Information
By construction, $H_{k}=H(\tilde{Y}^{(k)})$ and $h_{k}=H(\tilde{Y}^{(k)}\mid X^{(k)})$, so $H_{k}-h_{k}=I(X^{(k)};\tilde{Y}^{(k)})$. Mutual information is symmetric:
$I(A;B)=H(A)-H(A\mid B)=H(B)-H(B\mid A).$
Applying this with $A=X^{(k)}$ and $B=\tilde{Y}^{(k)}$ yields the stated identity. ∎
@@
$H_{k}-h_{k}$ is the expected information gain from conditioning on a specific context rather than using the averaged predictive distribution. The symmetric form shows it is equally the expected information gained about the sampled context from observing a token drawn from the model-induced next-token distribution.
@@ 5.6 Theoretical Analysis of the Information-Gain Span
$\operatorname{IGS}(k_{\mathrm{s}},k_{\mathrm{l}}):=u_{k_{\mathrm{s}}}\bigl(1-u_{k_{\mathrm{l}}}\bigr),\qquad u_{k}:=h_{k}/H_{k}.$
Theorem 3 (IGS and Markov order (regime separation)).
Assume the text source is an ergodic Markov chain of exact order $m$ and the model predicts perfectly. Assume further that $H_{k}$ is non-increasing. Then $u_{k}$ is non-increasing for $k<m$ and non-decreasing for $k\geq m$ (it may be flat for $k\geq m$). In particular, $u_{k}$ attains a minimum at $k=m$, though the minimiser need not be unique.
@@
Step 1: Behaviour of $u_{k}$. For $k<m$, $h_{k}$ falls strictly and $H_{k}$ does not rise, so $u_{k}$ falls. For $k\geq m$, $h_{k}$ is constant ($h_{m}$) while $H_{k}$ is non-increasing, so $u_{k}=h_{m}/H_{k}$ is non-decreasing or stays flat. Thus $u_{k}$ attains a minimum at $k=m$ (the minimiser need not be unique if $H_{k}$ is flat for $k\geq m$).
Step 2: The factor $(1-u_{k_{\mathrm{l}}})$. This term is largest when $u_{k_{\mathrm{l}}}$ is smallest. Under the theorem’s conclusion, $u_{k}$ attains its minimum at $k=m$ (and may remain flat on an interval). Hence the maximal value of $(1-u_{k_{\mathrm{l}}})$ is achieved by choosing $k_{\mathrm{l}}$ at a minimiser of $u_{k}$, i.e. $k_{\mathrm{l}}=m$ (or any $k_{\mathrm{l}}\geq m$ if $u_{k}$ is flat for $k\geq m$). In practice, one may still choose a larger $k_{\mathrm{l}}$ to represent a “long-context” regime, even when this does not improve the idealised bound.
Step 3: The factor $u_{k_{\mathrm{s}}}$. Because $u_{k}$ falls on $k<m$, $u_{k_{\mathrm{s}}}$ is largest when $k_{\mathrm{s}}$ is as small as allowed, subject to $k_{\mathrm{s}}\leq m$. ∎
@@
Under a finite-memory source, $u_{k}$ decreases until the relevant history is captured (around $m$) and then ceases to improve. This motivates selecting $k_{\mathrm{s}}\leq m<k_{\mathrm{l}}$ so that IGS contrasts a pre-saturation short-range regime with a post-saturation long-range regime. Using IGS peaks to estimate an effective memory scale is therefore heuristic and would require additional assumptions for a formal guarantee.
@@ 6 Conclusion
We proposed a task-agnostic framework for probing how large language models process information by tracking predictive uncertainty as context grows. The core object is the Entropy Decay Curve (EDC), defined by plotting the length-conditional uncertainty index $u_{k}=h_{k}/H_{k}$ across context lengths $k$. Unlike single-number summaries, the EDC exposes a model’s full trajectory from short-context ambiguity to long-context stabilisation, yielding a quantitative cognitive profile.
Across three open models and three public-domain corpora, we observed stable and interpretable differences in these profiles. Larger models tended to show a sharper transition from high initial uncertainty to low long-context uncertainty, while smaller models often exhibited a slower decay and a higher plateau on more difficult texts. The same analysis revealed strong corpus effects: texts that become locally predictable with sufficient context produced markedly lower long-context uncertainty than stylistically or semantically demanding works.
To summarise a decay pattern with a single scalar, we introduced the Information Gain Span (IGS), which contrasts a short-context regime with a long-context regime. Empirically, IGS tracked the qualitative desirability of a profile—high initial uncertainty coupled with strong long-range resolution—and enabled direct comparison across model–corpus pairs.
Beyond comparison, EDCs also act as a diagnostic. In particular, near-zero long-context uncertainty (entropy collapse) is difficult to reconcile with genuine generalisation on natural text and therefore serves as a practical warning signal for memorisation or benchmark contamination.
Finally, we connected the estimators to population quantities under the same sliding-window sampling rule and provided sufficient conditions under which $h_{k}$ is monotone in $k$. We also clarified that monotonicity of $H_{k}$ is not guaranteed in general, and interpreted $H_{k}-h_{k}$ as an expected information gain under the induced context distribution.
Limitations and future work.
Our study is limited to a small set of models, corpora, and window lengths, and relies on finite-sample Monte Carlo estimates. Future work should (i) extend coverage to more architectures and training regimes, (ii) evaluate specialised domains (code, scientific, medical, legal), (iii) characterise estimator variance and sensitivity to sampling design, and (iv) refine contamination tests by combining EDC signals with independent memorisation audits.